Enough of a monoprotic acid is dissolved in water to produce a 0.0146 M solution. The pH of the resulting solution is 2.41. Calculate the Ka for the acid
Answer #1

Ka = [H+][A-]/[HA]

[H+] = 10^-pH = 10^-2.41 = 0.00389

[H+] = [A-] = 0.00389

[HA] = 0.0146-0.00389 = 0.01071

then

Ka = (0.00389*0.00389)/(0.01071) = 0.00141289